The Dapdap tree in the Philippines is prone to wasp infestation. Suppose that the percent P of defoliation is approximated by P= 300/3+17e^-0.0625xwhere x is the number of egg masses in thousands. If the percent of defoliation is 20%, approximately how many egg masses are there?

1
Expert's answer
2021-10-14T05:48:00-0400

"P=20 \\
20 = \frac{300}{3 +17e^{-0.0625x}} \\
2 = \frac{30}{3 +17e^{-0.0625x}} \\
15 = 3 +17e^{-0.0625x} \\
17e^{-0.0625x} = 12 \\
e^{-0.0625x} = \frac{12}{17} \\
-0.0625x = ln(\frac{12}{17}) \\
x = \frac{-0.348306}{-0.0625} \\
x = 5.572 ≈ 6"